How many atoms are in 15.0 moles of C2H6O?
1
Expert's answer
2016-04-21T11:50:04-0400
Na(the Avogadro constant) - is the number of particles(atoms, molecules…) in 1 mol. Na= 6,02*1023
ν – the number of mols
N = Na*ν = 6.02*1023 * 15.0 = 90.3*1023